OpenStudy (katie0146722):

Explain the steps you would use to compare and . Remember, you can use the forward-slash key (/) to type fractions. For example, type as

OpenStudy (katie0146722):

Explain the steps you would use to compare four-fifteenths and five-sixths. Remember, you can use the forward-slash key (/) to type fractions. For example, type one-fourth as one slash four.

OpenStudy (katie0146722):

@hartnn

OpenStudy (katie0146722):

@AakashSudhakar @ajn946 @angelicwolf @bedpotato @Batra @britt_nayyCx @bigryan305 @Callisto @cutegirl @ChasingMaggie @Chineseboy15 @dan815 @DeadShot @DeeBush96 @Dara08

OpenStudy (katie0146722):

please help

OpenStudy (anonymous):

All right.

OpenStudy (katie0146722):

thnx >.<

OpenStudy (katie0146722):

@Chineseboy15 r u still there?

OpenStudy (anonymous):

First, make the two fractions have same denominators. \[\frac{ 4\times4 }{ 15\times4 }, \frac{ 5\times10 }{ 6\times10 }\] Then we get \[\frac{ 16 }{ 60 }<\frac{ 50 }{ 60 }\].

OpenStudy (anonymous):

Yes, I have just finish my answer.

OpenStudy (katie0146722):

oh i understand now thnx

OpenStudy (anonymous):

You are welcome!